OpenStudy (anonymous):

The speed of a passenger train is 18 mph faster than the speed of a freight train. The passenger train travels 270 miles in the same time it takes the freight train to travel 180 miles. Find the speed of each train.

OpenStudy (anonymous):

Let the speed of the freight train be 'v mph' The speed of the passenger train is thus 'v+18 mph' v=d/t t=d/v Therefore, 270/(v+18)=180/v or, 270v=180v + 3240 or, v=36 mph Thus the speeds are 36 and 54 mph!